# Doping-Driven Modulation of Spin–Orbit Coupling, Spin Textures, and Rashba–Edelstein Response in Chiral Tellurium: A First-Principles Study

### v2

@@ Abstract
Chiral semiconductors such as elemental tellurium (Te) exhibit unconventional spin textures and large charge‐to‐spin conversion efficiencies, yet the influence of doping on these properties remains underexplored. Here, we address this gap by investigating how substituting Te with lighter (S, Se) or heavier (Sb) elements systematically modifies the spin‐orbit–driven phenomena in chiral Te, including the band structure, spin Berry curvature, and Rashba–Edelstein response. The objective is to determine whether doping strategies can be leveraged to optimize collinear spin textures, enhance spin accumulation, and possibly extend spin lifetimes—all crucial aspects for magnet‐free spintronics. Using density functional theory calculations implemented in Quantum ESPRESSO, combined with tight‐binding interpolation in PAOFLOW, we map out the doping‐dependent electronic states and quantify their associated spin transport coefficients. Our findings reveal that lighter dopants shift the Fermi level to regions of pronounced spin splitting, thereby increasing the magnitude of spin‐current conversion, whereas heavier dopants can introduce or remove near‐degenerate bands that strongly affect spin‐orbit coupling. In both scenarios, the fundamental chirality of Te remains robust, preserving the radial or “collinear” spin‐momentum locking. These results not only confirm that doping is a potent and feasible route for tuning spin‐orbit phenomena but also offer practical guidelines for experimental efforts aiming to engineer chiral semiconductors for spin devices. By correlating dopant identity with specific spin‐texture enhancements, this study paves the way for rationally designing next‐generation spintronic components free from external magnetic fields.
@@ I Introduction
Despite this progress, many open questions remain regarding how to systematically tune or engineer the spin-orbit–mediated response in these chiral materials. One particular gap lies in understanding the role of doping and carrier concentration. Although doping is frequently mentioned in the broader field of spintronics as a method to shift Fermi levels or amplify spin splitting [24, 8, 25], a rigorous first-principles exploration of how doping strategies—either via rigid band shifts or substituting heavier (or lighter) atoms—affect the spin textures, spin relaxation times, and charge-to-spin conversion efficiencies remains incomplete [2, 3, 14, 15]. Indeed, doping elemental Te or a related chiral system offers a compelling way to enhance or suppress specific band-structure features (such as Weyl-like crossings) and thus to realize “persistent spin helix” conditions [17, 32, 38]. Such doping-based manipulation of spin-orbit–driven phenomena could be critical for high-efficiency, long-distance spin transport.
Here, we address this gap by presenting a systematic computational study of doping in a prototypical chiral semiconductor (elemental Te) to clarify how electron/hole injection, as well as heavier-element substitution, modify the spin texture and spin lifetime. While the references above have thoroughly examined pristine crystals, doping remains relatively unexplored, particularly from a quantitative ab initio perspective. By identifying specific doping levels or substitutional approaches that maximize charge-to-spin conversion and promote stable, long-lived spin accumulation, this work aims to offer a practical design strategy for chiral spintronics devices. Importantly, the modest computational cost of elemental Te—along with well-developed workflows for doping and spin-orbit calculations—makes this project both feasible on a single laptop and poised for high impact, as optimizing doping provides a direct knob to tune spin-orbit phenomena in chiral materials [36, 18].
@@ II Methodology
In this study, all first-principles calculations were performed within the density functional theory (DFT) framework as implemented in the Quantum ESPRESSO (QE) suite [43]. We employed fully relativistic pseudopotentials that incorporate spin-orbit coupling (SOC) in order to correctly capture the relativistic effects relevant for chiral semiconductors. The Perdew–Burke–Ernzerhof (PBE) exchange-correlation functional [44] was chosen because it offers a good balance between accuracy and computational cost for main-group semiconductors such as elemental Te. Convergence tests were carried out to establish suitable kinetic energy cutoffs for the plane-wave basis sets, as well as appropriately dense k-point meshes (Monkhorst–Pack grids) sufficient to resolve the subtle spin splitting phenomena near the Fermi level. Whenever a small supercell was introduced to simulate doping (by substituting one Te site with a heavier or lighter element), the structure was fully relaxed using the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm until the residual forces and total energies were below stringent convergence thresholds. This step ensured that any local distortion or relaxation caused by the dopant would be properly accounted for while preserving the overall chirality of the system.
After obtaining self-consistent charge densities and band structures in Quantum ESPRESSO, we used the PAOFLOW code [20, 39] for post-processing analyses of spin textures, Berry curvatures, and derived spin transport coefficients. PAOFLOW constructs a tight-binding representation in a projected atomic-orbital basis, allowing fine interpolation of the bands throughout the Brillouin zone. This interpolation is particularly advantageous for computing spin-texture maps, Rashba–Edelstein coefficients, and spin Hall conductivities, all of which rely on dense sampling in momentum space. For each doping configuration (electron- or hole-like doping by chemical substitution), the Fermi level shift was examined to determine changes in the low-energy electronic structure. Subsequently, spin and orbital polarizations were tracked as functions of wavevector and energy, enabling a direct comparison between pristine and doped cases. Where possible, we benchmarked the pristine Te results against literature data [25] to confirm the reliability of our computational setup.
@@ III.1 Band Structure
The band structures obtained for pristine tellurium (Te) and its three doped variants (Te:S, Te:Se, and Te:Sb) reveal several noteworthy trends that shed light on how substitutional doping alters the electronic states near the Fermi level. In the pristine Te band plot, the valence and conduction bands show characteristic dispersions associated with chiral Te, including regions of relatively narrow bandwidth near the Fermi level (set to 0 eV for plotting). These results are consistent with previous reports that have identified Te as a narrow-gap semiconductor whose chiral symmetry can lead to unusual spin textures and spin-momentum lockings [32, 25]. The overall gap between the highest occupied and lowest unoccupied states in our pristine Te calculations lies near 0 eV in the figure because of the chosen alignment, but the relative separation between valence and conduction bands and the presence of avoided crossings in certain Brillouin zone directions are qualitatively in line with known studies of elemental Te.
@@
Upon substituting Te by sulfur (Te doped with S), the band structure shows a modest upward shift of states near the Fermi energy, suggesting that lighter-element doping in Te can modify the internal chemical pressure and lattice parameters, thereby altering the relative positions of the valence and conduction bands. This effect is often referred to as “chemical pressure doping” and has been reported in other chalcogenide systems as well [25]. In our calculations, the energy dispersions appear slightly more spread in the conduction region for Te:S than in pristine Te, pointing to a change in effective masses of carriers. Similar trends of band-edge shifts induced by isoelectronic but lighter dopants have been observed experimentally in various chiral and non-centrosymmetric semiconductors [32].
@@
Doping with selenium (Te doped with Se), another group 16 element closer in atomic number to Te, produces a more moderate perturbation in the band structure. Compared to the Te:S case, the band edges in Te:Se remain somewhat closer to those of pristine Te, consistent with the lesser size mismatch and more similar chemical properties between Se and Te. The conduction bands display a slightly narrower dispersion around the Fermi level, which can be attributed to subtle lattice relaxations and spin-orbit coupling (SOC) changes introduced by Se. These observations align with prior studies showing that Se substitution in Te can tune electronic and spin-orbit–driven properties without drastically altering the overall semiconducting nature [33].
@@
The most pronounced alteration is observed when Te is doped with antimony (Te doped with Sb). Here, the band structure reveals a set of bands crossing or sitting very near 0 eV, indicative of a partially metallic or heavily doped character. Since Sb belongs to group 15, its substitution can effectively change the electron count in the lattice, often leading to hole-type doping in tellurium-based systems [46]. In our band plots, the Sb-doped sample shows states that either cross or pin to the Fermi energy more strongly than in the other doped variants, suggesting that Sb doping may offer a route to achieving p-type conduction with significant changes to the density of states at $E_{F}$. These findings corroborate earlier theoretical and experimental works demonstrating that heavier group 15 dopants can shift the Fermi level into regions of higher density of states, thus altering transport and spin properties [32].
@@
Taken together, these results demonstrate that chiral elemental Te is highly sensitive to substitutional doping, with relatively modest atomic replacements leading to notable shifts or redistributions of the bands near the Fermi level. From the perspective of engineering spin-orbit–related responses—such as radial spin textures or Rashba–Edelstein effects—these band shifts are significant. In the context of existing literature, our findings align with the suggestion that doping provides a powerful handle to tune the magnitude and location of spin-split states, and hence to optimize spin-charge conversion phenomena [25, 32]. Specifically, lighter dopants (S, Se) mostly cause moderate shifts and slight modifications of the dispersion, whereas Sb doping can drastically move the Fermi level into or out of particular bands, potentially enhancing or suppressing spin-polarized currents. These doping-induced modifications pave the way for future calculations of spin textures, Berry curvatures, and spin relaxation times, complementing the pioneering reports on collinear spin transport and large charge-to-spin conversion in chiral semiconductors [33, 49].
@@ III.2 Spin Berry Curvature
The spin Berry curvature plots for pristine Te and the three doped variants (Te:S, Te:Se, Te:Sb) exhibit pronounced changes in both the magnitude and sign of the curvature at various high-symmetry points in the Brillouin zone. For pristine Te, the curvature shows multiple oscillations between positive and negative values along the high-symmetry paths, with notable peaks around the L and H points. Such oscillatory behavior is consistent with previous reports of a radial or collinear spin texture in elemental Te, where the interplay of chiral symmetry and spin-orbit coupling (SOC) leads to sizable Berry curvature contributions near certain band crossings [25]. The highest peak observed near the L–H region in pristine Te underlines the strong SOC effects intrinsic to Te’s chiral crystal structure, a factor that has been linked to the large spin-splitting phenomena and potential long-range spin transport in this material.
@@
Upon substitutional doping with S, the overall amplitude of the spin Berry curvature is notably reduced except for a few spikes near $\Gamma$ and H. This reduction can be attributed to the lighter atomic mass of sulfur compared to tellurium, which tends to lower the effective SOC strength. Consequently, while there are still visible peaks, especially near A and H, the diminished curvature suggests a shift in the electronic bands that smears out or weakens some of the sharp features observed in pristine Te. These findings align with earlier theoretical predictions that lighter chalcogen dopants can quench some of the SOC-driven band splittings, thereby modifying the spin-momentum locking and the resultant Berry curvature [45].
@@
Doping with Se, on the other hand, produces one particularly large and sharp peak of the spin Berry curvature, especially around the L point, exceeding values of 500 Å2. Because selenium is chemically similar to tellurium but has a slightly lower atomic number, this behavior indicates that the doping introduces localized band inversions or anticrossings that can drastically enhance the curvature at specific k-points. Such strong, localized peaks are often indicative of near-degenerate bands that are split by SOC in a chiral crystal, consistent with previous observations in closely related chiral materials where moderate changes in atomic potential lead to pronounced Berry curvature “hot spots” [43]. The sharpness of the feature further suggests that Se doping may position the Fermi level in a regime of near-crossings, thereby magnifying the intrinsic contributions to spin-dependent transport.
@@
Finally, doping with Sb reveals both positive and negative excursions of the spin Berry curvature, with a conspicuously deep negative dip near A and a sizable positive peak beyond 200 Å2 near L. Because antimony is heavier than sulfur or selenium, its presence tends to enhance SOC effects overall. However, it can also introduce additional carriers (depending on the valence state) and potentially shift the Fermi level into different bands compared to pristine Te. The simultaneous appearance of large negative and positive values underlines a more intricate band restructuring, where one set of states might reinforce the curvature whereas another set counters it, leading to sign changes that are more pronounced than in the S- or Se-doped cases. Such counteracting contributions are reminiscent of doping-driven band inversions observed in other topologically nontrivial systems, wherein heavy-element substitution can open gaps and generate additional Berry curvature pockets in momentum space [47, 48].
@@
These doping-induced variations confirm that controlled elemental substitution can selectively tune the bands where SOC is most active, thus reshaping the spin Berry curvature profile. Prior experimental and theoretical work on chiral semiconductors has emphasized the importance of locating “hot spots” of Berry curvature and spin-momentum locking near the Fermi level for possible spintronic applications [25]. Our results show that S and Se doping can either weaken or localize strong curvature features, while Sb doping can induce both enhancement and sign inversion. These observations indicate that doping not only modifies the magnitude of the overall spin-related response but can also reorganize the band structure in ways that may facilitate or hinder long-range spin transport. From a device-design standpoint, the possibility of achieving large positive or negative Berry curvature at specific k-points suggests clear strategies for optimizing chiral Te for magnet-free spintronics: lighter dopants (like S) may reduce extrinsic scattering while heavier dopants (like Sb) may amplify SOC-driven phenomena. Hence, identifying the dopant that maximizes desirable curvature features provides a pathway to engineering higher charge-to-spin conversion efficiencies and more robust spin accumulation, directly addressing the open questions about doping strategies raised in previous chiral spintronics studies.

### v3

# Modulation of Spin–Orbit Coupling, Spin Textures, and Rashba–Edelstein Response in Chiral Tellurium: A First-Principles Study

@@ Abstract
Chiral semiconductors such as elemental tellurium (Te) exhibit unconventional spin textures and large charge‐to‐spin conversion efficiencies, yet the influence of introducing elements on these properties remains underexplored. Here, we address this gap by investigating how substituting Te with lighter (S, Se) or heavier (Sb) elements systematically modifies the spin‐orbit–driven phenomena in chiral Te, including the band structure, spin Berry curvature, and Rashba–Edelstein response. The objective is to determine whether elemental substitution strategies can be leveraged to optimize collinear spin textures, enhance spin accumulation, and possibly extend spin lifetimes—all crucial aspects for magnet‐free spintronics. Using density functional theory calculations implemented in Quantum ESPRESSO, combined with tight‐binding interpolation in PAOFLOW, we map out the element‐dependent electronic states and quantify their associated spin transport coefficients. Our findings reveal that lighter elements shift the Fermi level to regions of pronounced spin splitting, thereby increasing the magnitude of spin‐current conversion, whereas heavier elements can introduce or remove near‐degenerate bands that strongly affect spin‐orbit coupling. In both scenarios, the fundamental chirality of Te remains robust, preserving the radial or “collinear” spin‐momentum locking. These results not only confirm that introducing elements is a potent and feasible route for tuning spin‐orbit phenomena but also offer practical guidelines for experimental efforts aiming to engineer chiral semiconductors for spin devices. By correlating element identity with specific spin‐texture enhancements, this study paves the way for rationally designing next‐generation spintronic components free from external magnetic fields.
@@ I Introduction
Despite this progress, many open questions remain regarding how to systematically tune or engineer the spin-orbit–mediated response in these chiral materials. One particular gap lies in understanding the role of elemental substitution and carrier concentration. Although introducing elements is frequently mentioned in the broader field of spintronics as a method to shift Fermi levels or amplify spin splitting [24, 8, 25], a rigorous first-principles exploration of how elemental substitution strategies—either via rigid band shifts or substituting heavier (or lighter) atoms—affect the spin textures, spin relaxation times, and charge-to-spin conversion efficiencies remains incomplete [2, 3, 14, 15]. Indeed, introducing elements to Te or a related chiral system offers a compelling way to enhance or suppress specific band-structure features (such as Weyl-like crossings) and thus to realize “persistent spin helix” conditions [17, 32, 38]. Such elemental-based manipulation of spin-orbit–driven phenomena could be critical for high-efficiency, long-distance spin transport.
Here, we address this gap by presenting a systematic computational study of elemental substitution in a prototypical chiral semiconductor (elemental Te) to clarify how electron/hole injection, as well as heavier-element substitution, modify the spin texture and spin lifetime. While the references above have thoroughly examined pristine crystals, elemental substitution remains relatively unexplored, particularly from a quantitative ab initio perspective. By identifying specific substitutional approaches that maximize charge-to-spin conversion and promote stable, long-lived spin accumulation, this work aims to offer a practical design strategy for chiral spintronics devices [36, 18].
@@ II Methodology
In this study, all first-principles calculations were performed within the density functional theory (DFT) framework as implemented in the Quantum ESPRESSO (QE) suite [43]. We employed fully relativistic pseudopotentials that incorporate spin-orbit coupling (SOC) in order to correctly capture the relativistic effects relevant for chiral semiconductors. The Perdew–Burke–Ernzerhof (PBE) exchange-correlation functional [44] was chosen because it offers a good balance between accuracy and computational cost for main-group semiconductors such as elemental Te. Convergence tests were carried out to establish suitable kinetic energy cutoffs for the plane-wave basis sets, as well as appropriately dense k-point meshes (Monkhorst–Pack grids) sufficient to resolve the subtle spin splitting phenomena near the Fermi level. Whenever a element was introduced to Te crystal (by substituting one Te site with a heavier or lighter element), the structure was fully relaxed using the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm until the residual forces and total energies were below stringent convergence thresholds. This step ensured that any local distortion or relaxation caused by the dopant would be properly accounted for while preserving the overall chirality of the system.
After obtaining self-consistent charge densities and band structures in Quantum ESPRESSO, we used the PAOFLOW code [20, 39] for post-processing analyses of spin textures, Berry curvatures, and derived spin transport coefficients. PAOFLOW constructs a tight-binding representation in a projected atomic-orbital basis, allowing fine interpolation of the bands throughout the Brillouin zone. This interpolation is particularly advantageous for computing spin-texture maps, Rashba–Edelstein coefficients, and spin Hall conductivities, all of which rely on dense sampling in momentum space. For each doping configuration (electron- or hole-like doping by chemical substitution), the Fermi level shift was examined to determine changes in the low-energy electronic structure. Subsequently, spin and orbital polarizations were tracked as functions of wavevector and energy, enabling a direct comparison between pristine and element substitution cases. Where possible, we benchmarked the pristine Te results against literature data [25] to confirm the reliability of our computational setup.
@@ III.1 Band Structure
The band structures obtained for pristine tellurium (Te) and its three element substitution variants (Te2S, Te2Se, and Te2Sb) reveal several noteworthy trends that shed light on how substitutional element alters the electronic states near the Fermi level. In the pristine Te band plot, the valence and conduction bands show characteristic dispersions associated with chiral Te, including regions of relatively narrow bandwidth near the Fermi level (set to 0 eV for plotting). These results are consistent with previous reports that have identified Te as a narrow-gap semiconductor whose chiral symmetry can lead to unusual spin textures and spin-momentum lockings [32, 25]. The overall gap between the highest occupied and lowest unoccupied states in our pristine Te calculations lies near 0 eV in the figure because of the chosen alignment, but the relative separation between valence and conduction bands and the presence of avoided crossings in certain Brillouin zone directions are qualitatively in line with known studies of elemental Te.
@@
Upon substituting Te by sulfur, the band structure shows a modest upward shift of states near the Fermi energy, suggesting that lighter-element in Te can modify the internal chemical pressure and lattice parameters, thereby altering the relative positions of the valence and conduction bands. This effect has been reported in other chalcogenide systems as well [25]. In our calculations, the energy dispersions appear slightly more spread in the conduction region for Te2S than in pristine Te, pointing to a change in effective masses of carriers. Similar trends of band-edge shifts induced by isoelectronic but lighter elements have been observed experimentally in various chiral and non-centrosymmetric semiconductors [32].
@@
Substitution with selenium, another group 16 element closer in atomic number to Te, produces a more moderate perturbation in the band structure. Compared to the Te2S case, the band edges in Te:Se remain somewhat closer to those of pristine Te, consistent with the lesser size mismatch and more similar chemical properties between Se and Te. The conduction bands display a slightly narrower dispersion around the Fermi level, which can be attributed to subtle lattice relaxations and spin-orbit coupling (SOC) changes introduced by Se. These observations align with prior studies showing that Se substitution in Te can tune electronic and spin-orbit–driven properties without drastically altering the overall semiconducting nature [33].
@@
The most pronounced alteration is observed when Te is substituted with antimony. Here, the band structure reveals a set of bands crossing or sitting very near 0 eV, indicative of a partially metallic character. Since Sb belongs to group 15, its substitution can effectively change the electron count in the lattice, often leading to hole-type element in tellurium-based systems [46]. In our band plots, the sample shows states that either cross or pin to the Fermi energy more strongly than in the other element variants, suggesting that Sb may offer a route to achieving p-type conduction with significant changes to the density of states at $E_{F}$. These findings corroborate earlier theoretical and experimental works demonstrating that heavier group 15 elements can shift the Fermi level into regions of higher density of states, thus altering transport and spin properties [32].
@@
Taken together, these results demonstrate that chiral elemental Te is highly sensitive to substitutional element, with relatively modest atomic replacements leading to notable shifts or redistributions of the bands near the Fermi level. From the perspective of engineering spin-orbit–related responses—such as radial spin textures or Rashba–Edelstein effects—these band shifts are significant. In the context of existing literature, our findings align with the suggestion that element substitution provides a powerful handle to tune the magnitude and location of spin-split states, and hence to optimize spin-charge conversion phenomena [25, 32]. Specifically, lighter elements (S, Se) mostly cause moderate shifts and slight modifications of the dispersion, whereas Sb can drastically move the Fermi level into or out of particular bands, potentially enhancing or suppressing spin-polarized currents. These element-induced modifications pave the way for future calculations of spin textures, Berry curvatures, and spin relaxation times, complementing the pioneering reports on collinear spin transport and large charge-to-spin conversion in chiral semiconductors [33, 49].
@@ III.2 Spin Berry Curvature
The spin Berry curvature plots for pristine Te and the three elemental variants (Te2S, Te2Se, Te2Sb) exhibit pronounced changes in both the magnitude and sign of the curvature at various high-symmetry points in the Brillouin zone. For pristine Te, the curvature shows multiple oscillations between positive and negative values along the high-symmetry paths, with notable peaks around the L and H points. Such oscillatory behavior is consistent with previous reports of a radial or collinear spin texture in elemental Te, where the interplay of chiral symmetry and spin-orbit coupling (SOC) leads to sizable Berry curvature contributions near certain band crossings [25]. The highest peak observed near the L–H region in pristine Te underlines the strong SOC effects intrinsic to Te’s chiral crystal structure, a factor that has been linked to the large spin-splitting phenomena and potential long-range spin transport in this material.
@@
Upon substituting with S, the overall amplitude of the spin Berry curvature is notably reduced except for a few spikes near $\Gamma$ and H. This reduction can be attributed to the lighter atomic mass of sulfur compared to tellurium, which tends to lower the effective SOC strength. Consequently, while there are still visible peaks, especially near A and H, the diminished curvature suggests a shift in the electronic bands that smears out or weakens some of the sharp features observed in pristine Te. These findings align with earlier theoretical predictions that lighter chalcogen dopants can quench some of the SOC-driven band splittings, thereby modifying the spin-momentum locking and the resultant Berry curvature [45].
@@
With Se, on the other hand, produces one particularly large and sharp peak of the spin Berry curvature, especially around the L point, exceeding values of 500 Å2. Because selenium is chemically similar to tellurium but has a slightly lower atomic number, this behavior indicates that the element introduces localized band inversions or anticrossings that can drastically enhance the curvature at specific k-points. Such strong, localized peaks are often indicative of near-degenerate bands that are split by SOC in a chiral crystal, consistent with previous observations in closely related chiral materials where moderate changes in atomic potential lead to pronounced Berry curvature “hot spots” [43]. The sharpness of the feature further suggests that Se may position the Fermi level in a regime of near-crossings, thereby magnifying the intrinsic contributions to spin-dependent transport.
@@
Finally, substituting with Sb reveals both positive and negative excursions of the spin Berry curvature, with a conspicuously deep negative dip near A and a sizable positive peak beyond 200 Å2 near L. Because antimony is heavier than sulfur or selenium, its presence tends to enhance SOC effects overall. However, it can also introduce additional carriers (depending on the valence state) and potentially shift the Fermi level into different bands compared to pristine Te. The simultaneous appearance of large negative and positive values underlines a more intricate band restructuring, where one set of states might reinforce the curvature whereas another set counters it, leading to sign changes that are more pronounced than in the S- or Se-doped cases. Such counteracting contributions are reminiscent of elemental-driven band inversions observed in other topologically nontrivial systems, wherein heavy-element substitution can open gaps and generate additional Berry curvature pockets in momentum space [47, 48].
@@
These element-induced variations confirm that controlled elemental substitution can selectively tune the bands where SOC is most active, thus reshaping the spin Berry curvature profile. Prior experimental and theoretical work on chiral semiconductors has emphasized the importance of locating “hot spots” of Berry curvature and spin-momentum locking near the Fermi level for possible spintronic applications [25]. Our results show that S and Se can either weaken or localize strong curvature features, while Sb can induce both enhancement and sign inversion. These observations indicate that elemental substitution not only modifies the magnitude of the overall spin-related response but can also reorganize the band structure in ways that may facilitate or hinder long-range spin transport. From a device-design standpoint, the possibility of achieving large positive or negative Berry curvature at specific k-points suggests clear strategies for optimizing chiral Te for magnet-free spintronics: lighter elements (like S) may reduce extrinsic scattering while heavier elements (like Sb) may amplify SOC-driven phenomena. Hence, identifying the element that maximizes desirable curvature features provides a pathway to engineering higher charge-to-spin conversion efficiencies and more robust spin accumulation, directly addressing the open questions about doping strategies raised in previous chiral spintronics studies.